A spring scale reads 39.2 N when the object is hung vertically at rest. The object and spring scale are now swung in a vertical circle. Predict what the spring scale will read at the bottom of the circle if the object is moving at a speed of 6 m/s at that point and the radius of the circle is 1.200 m.
Answer #1

W = 39.2 N in = w - 39.2 = 3.9959 kg moz Tom( 3+x2) = 3.9959[3.81 4 4 2 3 = 159.0776 Newton